40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 9 13:40:02 2021, Ian MacMillan, Summary, Computer Scripts / Programs, SUS Plant Plan for New Optics 
    Reply  Tue Nov 9 16:55:52 2021, Ian MacMillan, Summary, Computer Scripts / Programs, SUS Plant Plan for New Optics IMG_8107.JPG
       Reply  Tue Nov 9 18:05:03 2021, Ian MacMillan, Summary, Computer Scripts / Programs, SUS Plant Plan for New Optics 
          Reply  Mon Nov 15 15:12:28 2021, Ian MacMillan, Summary, Computer Scripts / Programs, SUS Plant Plan for New Optics 
             Reply  Tue Nov 16 17:29:49 2021, Ian MacMillan, Summary, Computer Scripts / Programs, SUS Plant Plan for New Optics x_x_TF1.pdf
                Reply  Thu Nov 18 20:00:43 2021, Ian MacMillan, Summary, Computer Scripts / Programs, SUS Plant Plan for New Optics Final_Plant_Testing.zip
                   Reply  Mon Nov 22 16:38:26 2021, Tega, Summary, Computer Scripts / Programs, SUS Plant Plan for New Optics 
                      Reply  Mon Jan 24 17:10:25 2022, Tega, Summary, Computer Scripts / Programs, SUS Plant Plan for New Optics Screen_Shot_2022-01-24_at_5.33.15_PM.png
                         Reply  Thu Jan 27 16:40:57 2022, Tega, Summary, Computer Scripts / Programs, SUS Plant Plan for New Optics simplant_control_1.pngsimplant_control_0.png
                            Reply  Thu Feb 3 17:15:23 2022, Tega, Summary, Computer Scripts / Programs, SUS Plant Plan for New Optics Screen_Shot_2022-02-03_at_8.23.07_PM.png
                               Reply  Wed Feb 9 14:34:27 2022, Ian, Summary, Computer Scripts / Programs, SUS Plant Plan for New Optics SummaryMdemScreen.png
Message ID: 16626     Entry time: Thu Jan 27 16:40:57 2022     In reply to: 16615     Reply to this: 16645
Author: Tega 
Type: Summary 
Category: Computer Scripts / Programs 
Subject: SUS Plant Plan for New Optics 

[Ian, Paco, Tega]

Last night we set up the four main matrices that handle the conversion between the degrees of freedom bases and the sensor bases. We also wrote a bash script to automatically set up the system. The script sets the four change of bases matrices and activates the filters that control the plant. this script should fully set up the plant to its most basic form. The script also turns off all of the built-in noise generators. 

After this, we tried damping the optic. The easiest part of the system to damp is the side or y motion of the optic because it is separate from the other degrees of freedom in both of the bases. We were able to damp that easily. in attachment 1 you can see that the last graph in the ndscope screen the side motion of the optic is damped. Today we decided to revisit the problem.

Anyways, looking at the problem with fresh eyes today, I noticed the in pit2pit coupling has the largest swing of all the plant filters and thought this might be the reason why the inputs (UL,UR,LR,LL) to the controller was hitting the rails for pit DoF. I reduce the gain of the pit2pit filter then slowly increased it back to one. I also reduced the gain in the OSEM input filter from 1 to 1/100. The attached image (Attachment2) is the output from this trial. This did not solve the problem. The output when all OSEM input filter gain set to one is shown in Attachment2.

We will try to continue to tweak the coefficients. We are probably going to ask Anchal and Paco to sit down with us and really hone in on the right coefficients. They have more experience and should be able to really get the right values.

Attachment 1: simplant_control_1.png  198 kB  | Hide | Hide all
simplant_control_1.png
Attachment 2: simplant_control_0.png  197 kB  | Hide | Hide all
simplant_control_0.png
ELOG V3.1.3-